# Select, then Balance: A Plug-and-Play Framework for Exogenous-Aware Spatio-Temporal Forecasting

### v1

@@ C.1. Limitation
Second, our model currently assumes that spatio-temporal exogenous variables are numerical sensor signals. However, real-world scenarios often involve diverse, heterogeneous exogenous information, such as language and visual information. Recent research has explored using Large Language Models (LLMs) for time series (jin, 2024; Pan et al., 2024; Sun et al., [n. d.]; Jia et al., 2024; Chang et al., 2025; Liu et al., 2025b) and spatio-temporal forecasting (Liu et al., 2025a, 2024; Li et al., 2024; Xu and Liu, 2025), hoping that LLMs can gain numerical prediction capabilities through various alignment techniques. However, some studies have raised significant doubts about these approaches (Tan et al., 2024). Different from these views, we believe that language can be more effectively viewed as sparse, discrete exogenous auxiliary information. The key challenge lies in how to leverage this information to enhance various spatio-temporal tasks.
C.2. Future Direction
@@
For spatio-temporal reasoning tasks, we believe the core essence is to activate the intrinsic capabilities of language models to analyze the underlying causes of spatio-temporal changes and provide logical knowledge-based answers.

### v2

@@ D.1 Current Limitation
Second, our model currently assumes that spatio-temporal exogenous variables are numerical sensor signals. However, real-world scenarios often involve diverse, heterogeneous exogenous information, such as language and visual information (chen2024terra; luo2024towards; chen2025climateiqa). Recent research has explored using Large Language Models (LLMs) for time series (jin2024position; pan2024s; suntest; jia2024gpt4mts; chang2025llm4ts; liu2025timecma) and spatio-temporal forecasting (liu2024spatial; li2024urbangpt; liu2025st; xu2025gpt4tfp), hoping that LLMs can gain numerical prediction capabilities through various alignment techniques. However, some studies have raised significant doubts about these approaches (tan2024language). Different from these views, we believe that language can be more effectively viewed as sparse, discrete exogenous auxiliary information. The key challenge lies in how to leverage this information to enhance various ST tasks.
D.2 Future Direction
@@
For spatio-temporal reasoning tasks (han2025large; li2025stbench; hettige2025modular; ni2026streasoner), we believe the core essence is to activate the intrinsic capabilities of language models to analyze the underlying causes of ST changes and provide logical knowledge-based answers.